For Exercises 1-15, prove or disprove the given statement.


1. The product of any three consecutive integers is even.

2. The sum of any three consecutive integers is even.

3. The product of an integer and its square is even.

4. The sum of an integer and its cube is even.

5. Any positive integer can be written as the sum of the squares of two integers.

6. For a positive integer

Ꮧ Ꮧ +



7. For every prime number n, n + 4 is prime.

8. For n a positive integer, n > 2, n2- 1 is not prime.

9. For every positive integer n, n2 + n + 1 is prime.

10. For every positive integer n, 2n+ 1 is prime.

11. For n an even integer, n > 2, 2"1 - I is not prime.

12. The product of two rational numbers is rational.

13. The sum of two rational numbers is rational.

14. The product of two irrational numbers is irrational.

15. The sum of a rational number and an irrational number is irrational.
Answer #1

Page: integers The Ce ex: product of ny three consecutive enen 1, 2, 3 three consecubine integers 1X2X3=6 which is enen. Prod7201 Date: Page: egera 5. Any positive Integer can be written as the sum of the squares of two ex. x=130 13 can be written as- Date Page: 9. for eneng positive integer n n nt is not prime. ex: n = 4 ² n²+n+1 = 16+4+1 = 21 which is not a prime. I Dipp